# Riccardo Giordana Pozzi

### v1

@@ Results
The paper is organized as follows. Section 2 introduces the general structures of supersymmetric defect CFTs and sets the stage for constraining the order in which OPE coefficients enter the block expansion. In section 3 we discuss what are the assumptions under which a defect CFT may feature universal properties. We comment on the universal sectors when relaxing some of these assumptions and, finally, we provide the structure for the full NLO order contributions. We verify the universal property in a wide spectrum of supersymmetric defect CFTs realized by BPS Wilson lines in section 4 and we give a brief argument of universality from an holographic point of view in section 5. We conclude the work with a full analysis regarding the $1/2$ BPS Wilson line in Chern-Simons theories in section 6 and in section 7 we initiate the study of defect correlators of the $1/3$ BPS Wilson lines in ABJM at strong coupling. Appendix A contains further details on the ${\mathfrak{osp}}(2|4)$ superalgebra, the symmetry breaking pattern associated with the $1/2$ BPS line, and the orthogonality conditions entering the (super)block expansions.
@@ 2.1 Wilson lines as defect CFTs
where ${\mathcal{L}}(t)$ denotes a theory-dependent (super)connection [^1] In $3d$ theories, constructing $1/2$ BPS Wilson lines requires superconnections. For instance, in ABJM the $1/2$ BPS Wilson line is defined as the holonomy of a ${\mathfrak{u}}(N|N)$ superconnection [31].. The presence of such a line defect breaks the superconformal symmetry of the ambient theory, reducing it to a smaller subgroup. This residual symmetry generically includes, at least, the bosonic subalgebra ${\mathfrak{so}}(1,2)_{\text{conf}}\oplus{\mathfrak{so}}(d-1)_{\text{rot}}$, where $d$ is the spacetime dimension of the ambient Euclidean theory. In supersymmetric settings, only a part of the full supersymmetry algebra remains intact. Defects of this type are classified as $1/n$ BPS according to the fraction of supercharges they preserve.
Central features in the study of any CFT are the correlation functions. In the presence of a line defect, this naturally extends to the study of defect correlators, which involve local operator $\mathcal{O}_{i}(t_{i})$ inserted along the line, defined by
@@
In defect CFTs, a particularly distinguished set of defect operators are those related to the breaking of the bulk symmetries. A particular example that will be always present in the defect CFT is the displacement operator which is associated with the breaking of the translations orthogonal to the defect. When considering supersymmetric defects, the breaking pattern of supersymmetric currents leads to additional defect operators. Crucially, one can identify them by exploiting the Ward identity
@@
that explicitly relates broken generators to local insertions on the line. One can then further characterize the local operators by finding the representations under the preserved supergroup. By exploiting (super)Jacobi identities like
@@
one can explicitly read the charges of defect insertion $\mathbb{G}(t)$ by the algebra involving preserved $G$ and broken $\boldsymbol{G}$} generators
@@
Crucially, in these preliminary remarks we refrain from making explicit reference to symmetries, as their precise realization is theory-dependent. In specific models, symmetries naturally impose further constraints.
Our purpose here is instead to present a general and schematic description of the correlation functions of the displacement supermultiplet, particularly the four-point function. This is done in order to highlight the universal features common to all such defect theories. The following discussion should therefore be understood as theory-independent and schematic, to be refined when applied to concrete cases.
@@
Here, the $y_{i}$ denote the supersymmetric coordinates constructed from $t_{i}$, the coordinate along the line and $\theta_{i}$, the fermionic coordinates. The quantity $\langle ij\rangle$ is a supertranslation-invariant distance, whose precise form depends on the superspace, and thus on the specific theory, that the correlation function describes. All components except $t_{i}$ are theory-dependent, for an explicit example see subsection 6.2. The overall $C_{\mathcal{T}}$ is a physical theory-dependent constant related to the Bremsstrahlung function [9].
@@
Here, $\mathcal{Z}$ denotes a superconformal cross ratio. For chiral correlators, this is, for example, (6.6). More generally, the function can be expressed in terms of nilpotent invariants constructed from the Graßmann coordinates. Expanding in this basis yields a larger set of independent functions contributing to the correlators, see for example [20]. By expanding the left-hand side in superfield components and the right-hand side as a Taylor series in the fermionic coordinates, one can relate, term by term, the correlation functions of the multiplet components to derivatives of the superconformal invariant function.
@@
$z=\frac{t_{12}t_{34}}{t_{14}t_{32}}\in(-\infty,0)\,,\qquad\chi=\frac{t_{12}t_{34}}{t_{13}t_{23}}\in(0,1)$ (2.3)
@@
and with the mixed four-point with two displacement and two other generic operator in the multiplet
@@
Here $\mathbb{n}_{\mathbb{D}}$ and $\mathbb{n}_{\mathbb{D}\mathbb{O}}$ are theory-dependent constants that originate from the supersymmetric structure of the problem. While the two-point function of the superprimary is fixed solely in terms of ${\mathcal{C}}_{\mathcal{T}}$, with $\mathbb{n}_{\mathbb{P}}=1$, the inclusion of $Q$-descendants generally introduces additional constants, which arise from the action of the supercharges.
@@
for some theory dependent set of polynomials $\{p_{n}(z)\}$, see for example (6.8). When (2.2) has nilpotent invariants, one has also additional contributions coming from the higher-number of independent functions. These type of relations are exact and therefore hold at any order in perturbation theory. In particular, they remain valid when expanding around the generalized free field limit at strong coupling. In this regime, the leading behavior of the correlators is determined by the GFF, while higher-order corrections account for deviations.
@@
for some $\lambda_{\Delta}^{(0)}=\lambda_{\mathbb{D}\mathbb{D}\mathbb{T}_{\Delta}}^{(0)}$ representing the leading order three-point coefficients [^2] In the CPW expansion, the coefficients appearing are related to the three-point one by $\lambda_{\Delta}^{(0)}=k_{\mathbb{D}\mathbb{D}\mathbb{T}_{\Delta}}\lambda_{\mathbb{D}\mathbb{D}\mathbb{T}_{\Delta}}^{(0)}/(c_{\mathbb{D}\mathbb{D}}c_{\mathbb{T}\mathbb{T}})$, where $k_{\mathbb{D}\mathbb{D}\mathbb{T}_{\Delta}}$ denotes the normalization of the three-point function, and $c_{\mathbb{D}\mathbb{D}}$, $c_{\mathbb{T}\mathbb{T}}$ those of the two-point functions. However, for brevity, we neglect these normalization constant, specifying them explicitly only when necessary. of a generic conformal primaries $\mathbb{T}_{\Delta}$ having dimensions $\Delta$. In order to be exchanged we assume that it lies in the displacement OPE, i.e. $\mathbb{T}_{\Delta}\in\mathbb{D}\times\mathbb{D}$. The sum runs over all the conformal dimensions allowed by the OPE. Similarly one could do the same for (2.9), thus obtaining
@@
where $\gamma^{(1)}_{\Delta}$ are the anomalous dimension corresponding to the exchanged operators of dimension $\Delta$. In this first instance, we keep the order in $\epsilon_{\mathcal{T}}$ in the definitions of the NLO data. This leads to the identifications
@@
$\mathrm{g}_{\Delta}(z)=(-z)^{\Delta}\Bigl(\partial_{\tilde{\Delta}}g_{\tilde{\Delta}}(z)(-z)^{-\tilde{\Delta}}\Bigr)_{\tilde{\Delta}=\Delta}\,$,
@@ 2.2 Consistency constraints
In this subsection, we highlight some general consequences that follow from the identifications in (2.13) and (2.14). This line of reasoning will be systematically exploited in the next section to study the possible exchanged operators that can appear at NLO where, with additional assumptions, one can identify universal features. Before than this, however, we first derive constraints that can be obtained from more general considerations.
@@
can be interpreted as a constraint on the $\epsilon_{\mathcal{T}}$–order of $\lambda_{\Delta}^{(1)}$ and $\tilde{\lambda}_{\Delta}^{(1)}$. Crucially, as discussed above, the sum runs over all operator dimensions that may appear in the OPE. Operators that do not contribute at a given order have vanishing OPE coefficients, although they are still formally included in the sum.
@@
where by ${\mathcal{O}}(\epsilon_{\mathcal{T}})$ we mean that the order of the OPE can be either $\epsilon_{\mathcal{T}}$ or higher [^3] For clarity, with a slight abuse of notation these expression can be interpreted as $\lambda_{\Delta^{\star}}^{(1)}\tilde{\lambda}_{\Delta^{\star}}^{(0)}\gtrsim{\mathcal{O}}(\epsilon)$ and $\lambda_{\Delta^{\star}}^{(1)}\tilde{\lambda}_{\Delta^{\star}}^{(1)}\gtrsim{\mathcal{O}}(\epsilon)$ to stress that the order can be higher. This is equivalent to say that, for example, if a function is bounded by ${\mathcal{O}}(\epsilon^{2})$, then it is also bounded by ${\mathcal{O}}(\epsilon)$ contributions. Given that $\tilde{\lambda}_{\Delta}^{(0)}$ is constrained by the non-vanishing condition at GFF, $\tilde{\lambda}_{\Delta}^{(0)}\sim{\mathcal{O}}(1)$, consistency with the order in $\epsilon_{\mathcal{T}}$ of the left-hand-side leads to the requirement that either
@@
So, when considering the CPW expansion of correlators with external insertions of the same type, as in (2.13), the contributions associated with the $\mathbb{T}_{\Delta^{\star}}$ exchange given by
@@
will be either vanishing, if $\lambda^{(1)}_{\Delta^{\star}}=0$, or of higher $\epsilon_{\mathcal{T}}$-order, since $(\lambda^{(1)}_{\Delta^{\star}})^{2}\sim{\mathcal{O}}(\epsilon_{\mathcal{T}}^{2})$. In either case, one can conclude that $\mathbb{T}_{\Delta^{\star}}$ is not exchanged up to NLO.
@@
However, this may simply reflect an unfortunate choice of correlators. By instead considering correlators with different external insertions, one could further constrain the OPE coefficient through the procedure outlined above. On the other hand, if no such possibility arises, the $\epsilon_{\mathcal{T}}$-order obtained should be regarded as the correct one. For these cases, one may also note that if the $\Delta^{\star}$ is smaller than any $\Delta$ of the other exchanged operators, then the $\log(-z)$ contribution is absent. This implies that the exchanged operator must have vanishing anomalous dimension, thus pointing towards protected operators, see discussion in 3.2.1.
@@
Here, the sum runs over all $\Delta\in\mathrm{GFF}$, i.e. , over all scaling dimensions of the free theory. We are explicitly perturbing the leading-order data to NLO. The main difference with the previous analysis is that the sum does not include all operators that can formally appear in the OPE. The limitation of (2.19) arises when the spectrum of allowed operators changes at NLO. In such cases, to include all possible exchanges, $\Delta$ must be extended to run over all contributions allowed by the OPE, rather than being restricted to $\Delta\in\mathrm{GFF}$. This subtlety does not generate inconsistencies when the dimensions of the higher-order exchanges remain within the set of free-theory scaling dimensions. However, the situation changes when some of the operators appearing at NLO have classical dimensions outside the GFF set. If one does not account for this effect, the analysis that follows is restricted only to a partial sector.
@@
If the exchanged operators are only multiparticle states built from the external insertions, in the excluded contributions there would be also the exchange of the external field itself. Such a contribution, however, is only present when it is allowed by the OPE, e.g. $\mathbb{D}\in\mathbb{D}\times\mathbb{D}$ and it is related to the explicit breaking of the $\mathbb{Z}_{2}$ symmetry at NLO. At leading order the sum runs over $\{\Delta\in\mathbb{N}\mid\Delta\geq 2\Delta_{\mathbb{D}}\}$. Keeping this bound at NLO is equivalent to setting $\lambda_{\mathbb{D}\mathbb{D}\mathbb{D}}=0$ in (2.13). Importantly, as we discuss in 3.2.2, this coefficient controls all $\mathbb{Z}_{2}$-odd exchanges. Thus, setting it to zero automatically decouples these contributions, leaving the NLO analysis to consider only the $\mathbb{Z}_{2}$-even sector.

### v2

@@ Results
The paper is organized as follows. Sec. 2 introduces the general structures of supersymmetric defect CFTs and sets the stage for constraining the order in which OPE coefficients enter the block expansion. In Sec. 3 we discuss what are the assumptions under which a defect CFT may feature universal properties. We comment on the universal sectors when relaxing some of these assumptions and, finally, we provide the structure for the full NLO order contributions. We verify the universal property in a wide spectrum of supersymmetric defect CFTs realized by BPS Wilson lines in Sec. 4 and we give a brief argument of universality from an holographic point of view in Sec. 5. We conclude the work with a full analysis regarding the $1/2$ BPS Wilson line in Chern-Simons theories in Sec. 6 and in Sec. 7 we initiate the study of defect correlators of the $1/3$ BPS Wilson lines in ABJM at strong coupling. Appendix A contains further details on the ${\mathfrak{osp}}(2|4)$ superalgebra, the symmetry-breaking pattern associated with the $1/2$ BPS line, and the orthogonality conditions entering the (super)block expansions.
@@ 2.1 Wilson lines as defect CFTs
where ${\mathcal{L}}(t)$ denotes a theory-dependent (super)connection [^1] In $3d$ theories, constructing $1/2$ BPS Wilson lines requires superconnections. For instance, in ABJM the $1/2$ BPS Wilson line is defined as the holonomy of a ${\mathfrak{u}}(N|N)$ superconnection [30]. For other realizations, e.g. in ${\mathcal{N}}=4$, see [13, 29, 24] and $t=x^{0}\in(-\infty,\infty)$ the Euclidean time direction parametrizing an infinite straight line in $\mathbb{R}^{d}$. The presence of such a line defect breaks the superconformal symmetry of the ambient theory, reducing it to a smaller subgroup. This residual symmetry generically includes, at least, the bosonic subalgebra ${\mathfrak{so}}(1,2)_{\text{conf}}\oplus{\mathfrak{so}}(d-1)_{\text{rot}}$, where $d$ is the spacetime dimension of the ambient Euclidean theory. In supersymmetric settings, only a part of the full supersymmetry algebra remains intact. Defects of this type are classified as $1/n$ BPS according to the fraction of supercharges they preserve.
Central features in the study of any CFT are the correlation functions. In the presence of a line defect, this naturally extends to the study of defect correlators, which involve the local operators $\mathcal{O}_{i}(t_{i})$ inserted along the line, defined by
@@
In defect CFTs, a particularly distinguished set of defect operators are those related to the breaking of the bulk symmetries. When considering supersymmetric defects CFT, the breaking pattern of supersymmetric currents leads to additional operators. Crucially, one can identify them by exploiting the Ward identity
@@
that explicitly relates broken generators to local insertions on the line. We will refer to this set of defect operators obtained from Ward identities as fundamental (or elementary). One can then further characterize the local operators by finding the representations under the preserved supergroup. By exploiting (super-)Jacobi identities like
@@
one can explicitly read the charges of defect insertion $\mathbb{G}(t)$ by the algebra involving preserved $G$ and broken $\boldsymbol{G}$} generators.
A particular example that will be always present in the line defects is the displacement operator which is associated with the breaking of the translations orthogonal to the line. This is thus given by
$P^{i},{\mathcal{W}}]=\int{\mathcal{W}}[\mathbb{D}^{i}(t)]dt\,$, (2.2)
where the index runs up to the codimension of the defect i=1,,d1i=1,...,d-1, as expected from the ${\mathfrak{so}}(d-1)_{\text{rot}}$ [^2] For related discussions and a manifestly local realization of the displacement operator, see [15]..
@@
Crucially, in these preliminary remarks we refrain from making explicit reference to symmetries, as their precise realization is theory dependent. In specific models, symmetries naturally impose further constraints.
Our purpose here is instead to present a general and schematic description of the correlation functions of the displacement supermultiplet, particularly the four-point function. This is done in order to highlight the universal features common to all such defect theories. The following discussion should therefore be understood as theory independent and schematic, to be refined when applied to concrete cases.
@@
Here, the $y_{i}$ denote the supersymmetric coordinates constructed from $t_{i}$, the coordinate along the line and $\theta_{i}$, the fermionic coordinates. The quantity $\langle ij\rangle$ is a supertranslation-invariant distance, whose precise form depends on the superspace, and thus on the specific theory, that the correlation function describes. All components except $t_{i}$ are theory dependent; for an explicit example see Sec. 6.2. The overall $C_{\mathcal{T}}$ is a physical theory-dependent constant related to the bremsstrahlung function [15].
@@
Here, $\mathcal{Z}$ denotes a superconformal cross-ratio. For chiral correlators, this is, for example, (6.6). More generally, the function can be expressed in terms of nilpotent invariants constructed from the Graßmann coordinates. Expanding in this basis yields a larger set of independent functions contributing to the correlators; see for example [40]. By expanding the left-hand side in superfield components and the right-hand side as a Taylor series in the fermionic coordinates, one can relate, term by term, the correlation functions of the multiplet components to derivatives of the superconformal invariant function.
@@
$z=\frac{t_{12}t_{34}}{t_{14}t_{32}}\in(-\infty,0)\,,\qquad\chi=\frac{t_{12}t_{34}}{t_{13}t_{24}}\in(0,1)$ (2.4)
@@
and with the mixed four-point with two displacement and two other generic operators in the multiplet
@@
Here $\mathbb{n}_{\mathbb{D}}$ and $\mathbb{n}_{\mathbb{D}\mathbb{O}}$ are theory dependent constants that originate from the supersymmetric structure of the problem. While the two-point function of the superprimary is fixed solely in terms of ${\mathcal{C}}_{\mathcal{T}}$, with $\mathbb{n}_{\mathbb{P}}=1$, the inclusion of $Q$-descendants generally introduces additional constants, which arise from the action of the supercharges.
@@
for some theory-dependent set of polynomials $\{p_{n}(z)\}$; see for example (6.8). When (2.3) has nilpotent invariants, one has also additional contributions coming from the higher number of independent functions. These types of relations are exact and therefore hold at any order in perturbation theory. In particular, they remain valid when expanding around the generalized free-field limit at strong coupling. In this regime, the leading behavior of the correlators is determined by the GFF, while higher-order corrections account for deviations.
@@
for some $\lambda_{\Delta}^{(0)}=\lambda_{\mathbb{D}\mathbb{D}\mathbb{T}_{\Delta}}^{(0)}$ representing the leading order three-point coefficients [^3] In the CPW expansion, the coefficients appearing are related to the three-point one by $\lambda_{\Delta}^{(0)}=k_{\mathbb{D}\mathbb{D}\mathbb{T}_{\Delta}}\lambda_{\mathbb{D}\mathbb{D}\mathbb{T}_{\Delta}}^{(0)}/(c_{\mathbb{D}\mathbb{D}}c_{\mathbb{T}\mathbb{T}})$, where $k_{\mathbb{D}\mathbb{D}\mathbb{T}_{\Delta}}$ denotes the normalization of the three-point function, and $c_{\mathbb{D}\mathbb{D}}$, $c_{\mathbb{T}\mathbb{T}}$ those of the two-point functions. However, for brevity, we neglect these normalization constants, specifying them explicitly only when necessary. of a generic conformal primaries $\mathbb{T}_{\Delta}$ having dimensions $\Delta$. In order to be exchanged we assume that it lies in the displacement OPE, i.e. $\mathbb{T}_{\Delta}\in\mathbb{D}\times\mathbb{D}$. The sum runs over all the conformal dimensions allowed by the OPE. Similarly one could do the same for (2.10), thus obtaining
@@
where $\gamma^{(1)}_{\Delta}$ are the anomalous dimensions corresponding to the exchanged operators of dimension $\Delta$. In this first instance, we keep the order in $\epsilon_{\mathcal{T}}$ in the definitions of the NLO data. This leads to the identifications
@@
$\mathrm{g}_{\Delta}(z)=(-z)^{\Delta}\Bigl(\partial_{\tilde{\Delta}}g_{\tilde{\Delta}}(z)(-z)^{-\tilde{\Delta}}\Bigr)_{\tilde{\Delta}=\Delta}\,$, (2.14)
@@ 2.2 Consistency constraints
In this section, we highlight some general consequences that follow from the identifications in (2.15) and (2.16). This line of reasoning will be systematically exploited in the next section to study the possible exchanged operators that can appear at NLO where, with additional assumptions, one can identify universal features. Before this, however, we first derive constraints that can be obtained from more general considerations.
@@
can be interpreted as a constraint on the $\epsilon_{\mathcal{T}}$ order of $\lambda_{\Delta}^{(1)}$ and $\tilde{\lambda}_{\Delta}^{(1)}$. Crucially, as discussed above, the sum runs over all operator dimensions that may appear in the OPE. Operators that do not contribute at a given order have vanishing OPE coefficients, although they are still formally included in the sum.
@@
where by ${\mathcal{O}}(\epsilon_{\mathcal{T}})$ we mean that the order of the OPE can be either $\epsilon_{\mathcal{T}}$ or higher. [^4] For clarity, with a slight abuse of notation these expressions can be interpreted as $\lambda_{\Delta^{\star}}^{(1)}\tilde{\lambda}_{\Delta^{\star}}^{(0)}\gtrsim{\mathcal{O}}(\epsilon)$ and $\lambda_{\Delta^{\star}}^{(1)}\tilde{\lambda}_{\Delta^{\star}}^{(1)}\gtrsim{\mathcal{O}}(\epsilon)$ to stress that the order can be higher. This is equivalent to say that, for example, if a function is bounded by ${\mathcal{O}}(\epsilon^{2})$, then it is also bounded by ${\mathcal{O}}(\epsilon)$ contributions Given that $\tilde{\lambda}_{\Delta}^{(0)}$ is constrained by the nonvanishing condition at GFF, $\tilde{\lambda}_{\Delta}^{(0)}\sim{\mathcal{O}}(1)$, consistency with the order in $\epsilon_{\mathcal{T}}$ of the left-hand-side leads to the requirement that either
@@
So, when considering the CPW expansion of correlators with external insertions, of the same type, as in (2.15), the contributions associated with the $\mathbb{T}_{\Delta^{\star}}$ exchange given by
@@
will be either vanishing, if $\lambda^{(1)}_{\Delta^{\star}}=0$, or of higher $\epsilon_{\mathcal{T}}$ order, since $(\lambda^{(1)}_{\Delta^{\star}})^{2}\sim{\mathcal{O}}(\epsilon_{\mathcal{T}}^{2})$. In either case, one can conclude that $\mathbb{T}_{\Delta^{\star}}$ is not exchanged up to NLO.
@@
However, this may simply reflect an unfortunate choice of correlators. By instead considering correlators with different external insertions, one could further constrain the OPE coefficient through the procedure outlined above. On the other hand, if no such possibility arises, the $\epsilon_{\mathcal{T}}$ order obtained should be regarded as the correct one. For these cases, one may also note that if the $\Delta^{\star}$ is smaller than any $\Delta$ of the other exchanged operators, then the $\log(-z)$ contribution is absent. This implies that the exchanged operator must have vanishing anomalous dimension, thus pointing toward protected operators; see discussion in 3.2.1.
@@
Here, the sum runs over all $\Delta\in\mathrm{GFF}$, i.e. over all scaling dimensions of the free theory. We are explicitly perturbing the leading-order data to NLO. The main difference with the previous analysis is that the sum does not include all operators that can formally appear in the OPE. The limitation of (2.21) arises when the spectrum of allowed operators changes at NLO. In such cases, to include all possible exchanges, $\Delta$ must be extended to run over all contributions allowed by the OPE, rather than being restricted to $\Delta\in\mathrm{GFF}$. This subtlety does not generate inconsistencies when the dimensions of the higher-order exchanges remain within the set of free-theory scaling dimensions. However, the situation changes when some of the operators appearing at NLO have classical dimensions outside the GFF set. If one does not account for this effect, the analysis that follows is restricted only to a partial sector.
@@
If the exchanged operators are only multiparticle states built from the external insertions, in the excluded contributions there would be also the exchange of the external field itself. Such a contribution, however, is only present when it is allowed by the OPE, e.g. $\mathbb{D}\in\mathbb{D}\times\mathbb{D}$ and it is related to the explicit breaking of the $\mathbb{Z}_{2}$ symmetry at NLO. At leading order the sum runs over $\{\Delta\in\mathbb{N}\mid\Delta\geq 2\Delta_{\mathbb{D}}\}$. Keeping this bound at NLO is equivalent to setting $\lambda_{\mathbb{D}\mathbb{D}\mathbb{D}}=0$ in (2.15). Importantly, as we discuss in Sec. 3.2.2, this coefficient controls all $\mathbb{Z}_{2}$-odd exchanges. Thus, setting it to zero automatically decouples these contributions, leaving the NLO analysis to consider only the $\mathbb{Z}_{2}$-even sector.